OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function. y = (sin 3x)^lnx

OpenStudy (psymon):

Well, starting off, we just take the natural log of both sides so we can fix that power of ln(x) \[\ln (y) = \ln(x)*\ln(\sin3x)\]Since I took the natural log, I was able to bring down that ln(x) power and make it into a multiplication instead. So now the right side is a product rule derivative. Would you know how to do that differentiation?

OpenStudy (ness9630):

Tsk tsk, Psymon ofc.

OpenStudy (psymon):

Eh?

OpenStudy (anonymous):

1/y = (ln(3sin(3x))/(x) +3ln(x)cot(3x) so how do I solve for y?

OpenStudy (psymon):

You sure you have your derivative correct? \[f'(x)g(x) = f(x)g'(x)\] f'(x) = (1/x) g'(x) = 3cot(3x) \[\frac{ 1 }{ y }=\frac{ \ln(\sin(3x)) }{ x }+3\cot(3x)\]

OpenStudy (psymon):

Oops, meant for that to be a plus in the formula, not an equal sign of course.

OpenStudy (anonymous):

so now i have to solve for y right ?

OpenStudy (psymon):

Not really, no. When you do logarithmic differentiation, you're solving for dy/dx technically. What you do here is now multiply both sides by y to get: \[\frac{ dy }{ dx }=y(\frac{ \ln(\sin3x) }{ x }+3\cot3x)\]Now we replace y with the original problem. When we started, we said y = (sin3x)^lnx. Well, thats exactly what we substitute y for. So that gives us: \[\frac{ dy }{ dx }=(\sin3x)^{lnx}*(\frac{ \ln(\sin3x) }{ x }+3\cot3x)\] And that would be your answer. There is no real simplifying that much.

OpenStudy (anonymous):

Thanks for help .)

OpenStudy (psymon):

Yep. np :3